OpenStudy (loginusernamed):

Find the cube of 20.

OpenStudy (loginusernamed):

@ganeshie8 @gabbyalicorn @goalieboy @green_1

OpenStudy (loginusernamed):

@atreyu6s3x6

OpenStudy (anonymous):

the cube root of 20?

OpenStudy (anonymous):

like this?\[20^{3}\]

OpenStudy (anonymous):

or 20 cubed?

OpenStudy (loginusernamed):

40 B. 400 C. 800 D. 8000

OpenStudy (anonymous):

yea its cubed so 20*20*20 simple as that

OpenStudy (anonymous):

20*20*20=8000 (D)

OpenStudy (loginusernamed):

Thanks, BETS RESPONCE IS WHO GETS THIS RIGHT 1x1x1x1x1x2x22x1x43-122

OpenStudy (loginusernamed):

Jk

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

haha

OpenStudy (ddos_dragon):

824?

OpenStudy (ddos_dragon):

:3

OpenStudy (anonymous):

It's actually 1770...

OpenStudy (ddos_dragon):

You can literally take off all the ones. 2x22x43-122

OpenStudy (anonymous):

yea shes right xD

OpenStudy (loginusernamed):

Since ur still here, RANDOM POP QUESTION!!!

OpenStudy (ddos_dragon):

2 X 22 = 44 44 * 43 = 1892 1892 - 122 = 1770

OpenStudy (ddos_dragon):

I just put some random number in XD

OpenStudy (loginusernamed):

What is the volume of a cube with a side of length 14 units? A. 2744 cubic units B. 196 cubic units C. 84 cubic units D. 42 cubic units

OpenStudy (anonymous):

14*14*14=?

OpenStudy (ddos_dragon):

A?

OpenStudy (anonymous):

2744

OpenStudy (loginusernamed):

10*10*10=1000 then 4*4=8*4=32?

OpenStudy (anonymous):

4 * 4 = 16

OpenStudy (ddos_dragon):

4 X 16= 64

OpenStudy (anonymous):

wtf? xD

OpenStudy (loginusernamed):

How many 1 by 1 unit cubes would equal the volume of the figure shown? A. 1728 B. 224 C. 40 D. 24

OpenStudy (ddos_dragon):

B?

OpenStudy (anonymous):

Yes, B. 6*6*6=216 2*2*2=8 216+8= 224

OpenStudy (ddos_dragon):

2 X 2 X 2= 8 6 X 6 X 6= 216 216 + 8 = 224

OpenStudy (ddos_dragon):

You stole my thunder.

OpenStudy (loginusernamed):

so many ppl NOW RANDOM NON MATH QUESTION If bobby likes jane and jane is his ______ but jill is bobbys sister in law what is bobby to jane and vice versa

OpenStudy (loginusernamed):

What is the volume of the figure? A. 75 cubic units B. 125 cubic units C. 375 cubic units D. 750 cubic units

OpenStudy (anonymous):

5*5*5=125 125*3=375

OpenStudy (loginusernamed):

2. (1 pt) Find the volume of a cube with the given side length. 8 units A. 24 cubic units B. 64 cubic units C. 512 cubic units D. 4096 cubic units